OpenStudy (anonymous):

Find all the real cube roots of 0.000027 A. 0.0009 and -0.0009 B. 0.03 and -0.03 C. 0.0009 D. 0.03 Is the answer D?

OpenStudy (jadedry):

\[0.000027 = 27 *10^{-6}\] \[\sqrt[3]{27*10^{-6}} = 3*10^{-2}\] Your answer is correct! c:

OpenStudy (anonymous):

Thank you so much!! :) @Jadedry